OpenStudy (anonymous):

my question is which of these numbers are rational numbers but not integers -0.43 , -430 , 4.3 , -0.043 , 0.43 , -43 , 430 , 43 , 0.043 , -4.3 ?

OpenStudy (anonymous):

integers are whole numbers, and rational numbers are (basically) numbers that can be written as fractions. for example, .5 (1/2) is a rational number, but it is not an integer. can you do this problem from there?

OpenStudy (anonymous):

no ive tried and tried im just lost plz help ty

OpenStudy (anonymous):

well first eliminate the whole numbers, which are -430, -43, 430, and 43. that leaves you with: -0.43, 4.3, -0.043, 0.43, 0.043, -4.3 -0.43 = -43/100 4.3 = 43/10 -0.043 = -43/1000 .43 = 43/100 0.043 = 43/1000 -4.3 = -43/10 so... all of these except the whole numbers are rational, but not integers. (:

OpenStudy (anonymous):

ok now that i can see it like that it makes more sense . thank you so much

OpenStudy (anonymous):

you're welcome.